Project

General

Profile

11072013 Minutes
Attending: Dom, Susan, Nick, Himansu, Micheal, Joao, Kanika, Mat, Evan, Alex, Gavin

Operations: Run restarted on the swings. 6 dcms ran warm. Warm safe 350V on the new apds on diblock 3.

Ndos - Timing peak is back

Cold configuration is still hard to understand - Bring it up with Peter at Reco

Evan - Timing. Evan spoke with Neal Welcher. He though that the issue was with the ACC signal. Evan Rebooted the TDUs. Problems corrected themselves.
- Evan will check on the 3 sec prop delay events and make sure they at transmission errors.

Gavin - Numi is processing well.
11580-11590 (FTS does seem to know anything about these runs)
- Calibration is very green.

Dom - New on Watch Dog - New Table of runs.
- Gav asks for 1 month version
Tricia asks to divide out the partitions.
————————————————————————
Good runs discussion:
Nearline is not currently writing to the DB.
- Suggesting a offline data check.
- Database scheme is not optimal. Joao is working to build good runs from the onmon files.
- Kanika’s Ultra good list. Joao’s list 1136#.
- Joao will try to get the goods run list going again by next week.
————————————————————————

Slicer - Tomorrow.